OpenStudy (anonymous):

The reaction AB(aq)→A(g)+B(g) is second order in AB and has a rate constant of 0.0183L⋅mol−1⋅s−1 at 25.0 ∘C. A reaction vessel initially contains 250.0 mL of 0.200mol⋅L−1 AB which is allowed to react to form the gaseous product. The product is collected over water at 25.0 ∘C. How much time is required to produce 143.0mL of the products at a barometric pressure of 729.2mmHg . (The vapor pressure of water at this temperature is 23.8 mmHg.)

OpenStudy (anonymous):

@texaschic101

OpenStudy (aaronq):

First find the moles equivalent to that volume (143 mL). (i think you might need to subtract the VP of water from the total pressure). use: PV=nRT Find the moles left in the vessel with this information, then use the second order integrated rate law to find the time, t. \(\sf \dfrac{1}{[AB]_t} = kt + \dfrac{1}{[AB]_0}\rightarrow t=\dfrac{\dfrac{1}{[AB]_t} - \dfrac{1}{[AB]_0}}{k}\)

OpenStudy (anonymous):

I'm getting 1168s, does that sound about right?